OpenStudy (anonymous):

What is the mass of the air in a room measuring 10m x 5m x 2m. If the density of air is 1.3g/cm^3

TheSmartOne (thesmartone):

D=m/v

OpenStudy (anonymous):

So v= 10m x 5m x 2m = 100m^3 D = 1.3kg/m^3 D = M/V M= DxV m = 1.3kg/m^3 x 100m^3 M = 130 kg

TheSmartOne (thesmartone):

So 1.3=m/(10x5x2)

TheSmartOne (thesmartone):

So 1.3=m/100 m=130

TheSmartOne (thesmartone):

Wow OS took forever to post my reply.. -_-

OpenStudy (anonymous):

ikr I lost connection twice, but hey great minds think alike..

TheSmartOne (thesmartone):

:)

TheSmartOne (thesmartone):

@Matty13579

OpenStudy (anonymous):

@TheSmartOne But the density was 1.3g/cm^3 How did those change?

TheSmartOne (thesmartone):

demsity is 1.3 but the mass is 130kg

OpenStudy (anonymous):

@TheSmartOne What I don't understand is cm^3 x m^3